If a rectangle has length x and width x+13, what is the perimeter of rectangle in terms of x?

Answer by jim_thompson5910(35256)

We have two sides of x (they are opposite each other) and another pair of opposite sides of x+13. The opposite sides are equal in length.

There are four sides total:
side1 = x
side2 = x+13
side3 = x
side4 = x+13

The perimeter is the result of adding up all of the sides of the figure

perimeter = (side1)+(side2)+(side3)+(side4)
perimeter = (x)+(x+13)+(x)+(x+13)
perimeter = x+x+13+x+x+13
perimeter = (x+x+x+x)+(13+13)
perimeter = (4x)+(26)
perimeter = 4x+26